OpenStudy (anonymous):

A gymnast is trying to reach a velocity of 11 m/s at the end of a 16 meter runway. How fast must she accelerate? I came up with 28.54. Is this correct?

OpenStudy (anonymous):

Hmm that's rather off. Use the following kinematic equation for initial velocity \(v_i\), final velocity \(v_f\), acceleration \(a\), and displacement \(\Delta x\).\[v_f^2=v_i^2+2a\Delta x\]Hint: \(v_i\) here is implied to be zero.

OpenStudy (anonymous):

I think I'm a little confused as to what the displacement is. I set it up like 11 m/s= 0 + 2 (9.81) (16 m). Am I doing this completely wrong?

OpenStudy (anonymous):

The two things you got wrong: 1. You forgot to square the final velocity like in the formula and 2. you put in the acceleration of gravity for \(a\). Gravity has no role in this problem... in fact, \(a\) should not have anything substituted in for it at all, as it's what you're trying to find. Look at the statement you just wrote. It says 11=313.92, which clearly means that something was done wrong. However, you did get the displacement correct. :)

OpenStudy (anonymous):

i agree with yakeyglee the answer you should is 3.78125

OpenStudy (anonymous):

should have gotten*

OpenStudy (anonymous):

Okay, I think I finally have it figured out then. It's set up like 11^2= 0 + 2(16)?

OpenStudy (anonymous):

Where is \(a\)?